# Learning from Imperfect Demonstrations through Dynamics Evaluation

### 18 Dec 2023 v1

@@ IV-B Noise Angle Evaluation
As shown in Fig.4, we use a Better-Worse-Failed datasets (90 trajectories) to investigate the effect of a noise angle ranging from $10^{\circ}$ to $180^{\circ}$ on the success rate. The experimental results show that there exists an optimal noise angle $\theta_{noise}$ ranging from $20^{\circ}$ to $60^{\circ}$, which we refer to the optimal threshold.
It is important to note that when the noise angle is below the threshold, the algorithm’s performance improves as the noise angle increases. Conversely, when the noise angle is above the threshold, the performance declines as the noise angle increases. This phenomenon can be attributed to the fact that the threshold represents the true noise level in the data. Data above this threshold are real noise and has a negative impact on performance. The decrease of noise angle reduces the diversity of data, and when the noise angle is lower than the true noise level, it filters data that could potentially improve performance. Furthermore, algorithms that use our data confidence evaluation method can significantly improve performance at all angles and datasets, which highlights the strengths of our method. For instance, with the noise angle set to 30°, the success rate of CIQL(filter) is improved by $50\%$, CIQL-E by $53\%$, and CIQL-A by $60\%$ compared to the original algorithm..
(a) CIQL-A
(b) CIQL-E
Figure 5: Evaluation of Trajectories. We evaluate the trajectories using the reward function recovered by the optimal Q-function learned in the Better-Worse-Failed datasets (90, $\theta$${}^{\circ}$). The left side of the figure shows the results of CIQL-A and the right side shows the results of CIQL-E, where the red points are the Better data and the blue points are the Worse data.
@@ IV-C CIQL Evaluation
The horizon of a trajectory can indicate a human’s performance, specifically, the shorter horizon, the better its performance. We evaluate the human trajectories using the cumulative reward $\eta_{\tau_{i}}=\sum_{t=0}^{h_{i}}\gamma^{t}r(s_{t},a_{t})$. As shown in Fig. 5, the reward function inferred by CIQL-A aligns more closely with human intent. This is because CIQL-E simply filters noisy data, and noisy data does not contribute to the training of the reward function. On the other hand, CIQL-A’s reward function penalizes noisy data, leading to a reward function that more closely aligns human intent.

### 22 Feb 2024 v2

@@ IV-B Noise Angle Evaluation
We use a Better-Worse-Failed datasets (90 trajectories) to investigate the effect of a noise angle ranging from $10^{\circ}$ to $180^{\circ}$ on the success rate, as shown in Fig.4. The experimental results show that there exists an optimal noise angle $\theta_{n}$ ranging from $20^{\circ}$ to $60^{\circ}$, which we refer to the optimal threshold.
It is noteworthy that the performance of the algorithm exhibits a U-shaped relationship with the noise angle: it improves as the noise angle is below a certain threshold, yet it deteriorates as the noise angle exceeds this threshold. This behavior can be explained by the threshold’s role in delineating the true noise level within the data. Data points exceeding this threshold are considered real noise, which adversely affects performance. A decrease in the noise angle, while reducing data variability, may inadvertently filter out data that could potentially enhance performance.
Moreover, algorithms that employ our data confidence evaluation method demonstrate significant improvements in performance across a range of noise angles and datasets, illustrating the robustness and efficacy of our method. For example, when the noise angle is adjusted to 30°, we observe a significant enhancement in success rates: CIQL(filter) experiences a $50\%$ increase, CIQL-E shows a $53\%$ improvement, and CIQL-A achieves a $60\%$ boost in success rate relative to the baseline algorithm.
(a) Evaluate trajectories via CIQL-A
(b) Evaluate trajectories via CIQL-E
(c) Performance of CIQL-E(85.0%)
(d) Performance of CIQL-A(90.3%)
Figure 5: Performance of CIQL. We use the optimal policy and the Q-function learned in the Better-Worse-Failed datasets (90, $\theta$${}^{\circ}$) to test the performance of CIQL.
@@ IV-C CIQL Evaluation
The horizon of a trajectory can serve as a metric for human performance, with a reduced horizon often correlating with enhanced performance. We assess human trajectories through the cumulative reward, represented by the equation $\eta_{\tau_{i}}=\sum_{t=0}^{h_{i}}\gamma^{t}r(s_{t},a_{t})$. Figures 5(a) and 5(b) illustrate that the reward function recovered by CIQL-A aligns more closely with human intent. This is attributed to the fact that CIQL-E primarily filters out noisy data, which does not contribute to learn the reward function. Conversely, CIQL-A’s reward function actively penalizes such noisy data, resulting in a reward mechanism that is more in tune with human objectives. Moreover, as shown in Fig.5(c)and 5(c), even within the context of a successful outcome, CIQL-E exhibits certain limitations, notably a tendency to make contact with the desktop during execution. In contrast, CIQL-A demonstrates a superior ability to accurately secure the target object without any contact with the desktop throughout the task’s duration.